The height of a triangle is 3 inches less than 5 times the length of its base. The area is 34 square inches. Find the length of the base and the height of the triangle.
Found 2 solutions by rothauserc, MathTherapy:
Answer by rothauserc(4718)

area(A) of a triangle = (1/2) * base * height
:
let b be the triangle's base, then
:
height = 5b -3
:
(1/2) * b * (5b-3) = 84
:
multiply both sies of = by 2
:
b * (5b-3) = 168
:
5b^2 -3b = 168
:
5b^2 -3b -168 = 0
:
use the quadratic formula
:
x = (-b + or - square root(b^2 - 4 * a * c))/(2*a)
:
b = (-(-3) +square root((-3)^2 -4*5*(-168)))/(2*5) = 6.1043
:
b = (-(-3) -square root((-3)^2 -4*5*(-168)))/(2*5) = -5.5043
:
we reject the negative value for the base
:
*************************************************************
base is approximately 6.1 inches and height is 27.5 inches
:
(1/2) * 6.1 * 27.5 = 83.875 is approximately 84 square inches
:
answer checks
*************************************************************
:

Answer by MathTherapy(10810)


The height of a triangle is 3 inches less than 5 times the length of its base. The area is 34 square inches. Find the length of the base and the height of the triangle.
The other person's answer is WRONG, so ignore it!!
Correct answer: